# The Confidence Paradox: Can LLM Know When It’s Wrong?

### v1

@@ Abstract
Document Visual Question Answering (DocVQA) systems are increasingly deployed in real-world applications, yet they remain ethically opaque-often producing overconfident answers to ambiguous questions or failing to communicate uncertainty in a trustworthy manner. This misalignment between model confidence and actual knowledge poses significant risks, particularly in domains requiring ethical accountability. Existing approaches such as LayoutLMv3, UDOP, and DONUT have advanced state-of-the-art performance by focusing on architectural sophistication and accuracy; however, they fall short in ethical responsiveness.
To address these limitations, we introduce HonestVQA, a self-supervised honesty calibration framework for ethically aligned DocVQA. Our model-agnostic method quantifies uncertainty to identify knowledge gaps, aligns model confidence with actual correctness using weighted loss functions, and enforces ethical response behavior via contrastive learning. We further introduce two principled evaluation metrics-Honesty Score (H-Score) and Ethical Confidence Index (ECI)-to benchmark alignment between confidence, accuracy, and ethical communication. Empirically, HonestVQA improves DocVQA accuracy by up to 4.3% and F1 by 4.3% across SpDocVQA, InfographicsVQA, and SROIE datasets. It reduces overconfidence, lowering H-Score and ECI by 0.072 and 0.078, respectively. In cross-domain evaluation, it achieves up to 78.9% accuracy and 76.1% F1-score, demonstrating strong generalization. Ablation shows a 3.8% drop in accuracy without alignment or contrastive loss.
@@
Sahil Tripathi1, Md Tabrez Nafis2, Imran Hussain3, Jiechao Gao4thanks: Corresponding Author: jiechao@stanford.edu 1, 2, 3Jamia Hamdard, New Delhi, India 4Center for SDGC, Stanford University, California, USA
@@ 1 Introduction
Document Visual Question Answering (DocVQA) has emerged as a key challenge in multimodal AI Wang et al. (2025), enabling systems to answer questions based on visual and textual content in documents such as invoices, forms, contracts, and academic papers. These systems are widely deployed in enterprise automation Jiang et al. (2024), legal analysis Liu et al. (2024), and assistive technologies Zeng et al. (2024). However, despite their growing utility, DocVQA systems often lack ethical transparency—frequently returning confidently incorrect answers to ambiguous, adversarial, or under-specified queries. For instance, a system may assert the total invoice amount with high confidence even when the relevant table is partially occluded, or confidently misinterpret a scanned signature line as a date. Such failures can propagate serious downstream consequences, including legal misinterpretation, misinformation, or financial misjudgment.
However, the crux of the problem lies in the inability of existing DocVQA models to communicate uncertainty in a calibrated, ethically responsible manner. While State-of-the-Art (SOTA) systems such as LayoutLMv3111https://huggingface.co/microsoft/layoutlmv3-base Fujitake (2024), UDOP [^2] https://huggingface.co/microsoft/udop-large Wang et al. (2023a), and DONUT [^3] https://huggingface.co/naver-clova-ix/donut-base Li et al. (2024) focus on improving accuracy through sophisticated architecture and pretraining strategies, they fall short in aligning model confidence with actual knowledge. LayoutLMv3Fujitake (2024) tends to prioritize exact answers over conveying doubt, UDOP Wang et al. (2023a) frequently errs on the side of over-caution without actionable explanations, and DONUT Li et al. (2024) offers no uncertainty estimation at all—leading to ethically untrustworthy behavior in ambiguous scenarios. Therefore, recent advances in AI alignment research have emphasized the importance of ethical calibration Rao et al. (2023), including honesty Yang et al. (2024), confidence-awareness Stangel et al. (2025), and transparent failure modes Stewart et al. (2023). However, these insights have yet to be meaningfully integrated into DocVQA systems. To address these critical gaps, we propose HonestVQA, a self-supervised framework that calibrates model confidence to reflect its underlying knowledge and ethical responsibility. Our approach is model-agnostic and integrates three key components: (1) uncertainty quantification to identify knowledge gaps, (2) confidence-accuracy alignment through weighted loss optimization, and (3) contrastive learning to enforce ethical response boundaries in ambiguous contexts. We also introduce two novel evaluation metrics: i) Honesty Score (H-Score), which captures the alignment between confidence and correctness, and ii) Ethical Confidence Index (ECI), which evaluates whether high-confidence answers are ethically warranted.
@@ 2 Related Work
Recent research has increasingly focused on improving the reliability and interpretability of AI systems, especially in high-stakes domains. While core DocVQA models like have been discussed in Section 1, here we focus on complementary areas that our framework draws from—confidence calibration, ethical modeling, and contrastive learning. Confidence calibration techniques such as temperature scaling Xie et al. (2024) and label smoothing Müller et al. (2019) aim to align predicted probabilities with empirical accuracies. However, these methods are typically post-hoc and task-agnostic, often failing to generalize in multimodal settings. Selective prediction frameworks such as Chen et al. (2023) allow models to abstain from uncertain answers, but they usually rely on fixed thresholds and lack principled mechanisms to model epistemic uncertainty in visually grounded tasks like DocVQA. However, in the area of ethical and honest AI, efforts such as instruction tuning for alignment Zhang et al. (2023) and calibrated language modeling Zhu et al. (2023) emphasize epistemic humility—training models to express uncertainty when appropriate. However, these approaches are primarily developed for language-only models and remain underexplored in multimodal tasks involving structured visual data. Whereas, contrastive learning has shown strong performance in aligning multimodal representations, with frameworks like CLIP Gao et al. (2024) and ALIGN Wang et al. (2023b) leveraging contrastive objectives for image-text alignment. While effective at learning generalizable embeddings, such methods are not designed to enforce ethical boundaries or distinguish between honest and overconfident outputs in ambiguous scenarios.
@@
6:$\mathcal{L}_{\text{align}}\leftarrow\alpha\cdot\mathbb{1}[\hat{y}\neq y^{*}]\cdot\mathcal{C}+\beta\cdot\text{CE}(\hat{y},y^{*})$
@@
14:$\mathcal{L}_{\text{total}}\leftarrow\lambda_{1}\cdot\mathcal{L}_{\text{align}}+\lambda_{2}\cdot\mathcal{L}_{\text{contrast}}$
@@ 3 Methodology
As discussed earlier, HonestVQA is a model-agnostic calibration framework designed to enhance ethical transparency in DocVQA systems. It operates as a wrapper around pretrained DocVQA models (in our work, we evaluate our framework on top of pretrained models such as LayoutLMv3Fujitake (2024), UDOP Wang et al. (2023a), and DONUT Li et al. (2024) to demonstrate its generalizability), injecting uncertainty-aware alignment and contrastive reasoning to reduce overconfident yet incorrect outputs. The broader process of the HonestVQA is illustrated in Algorithm 1.
@@ 3.3 Contrastive Ethical Enforcement Module
To further refine the model’s response space under ambiguity, we introduce a contrastive loss that structurally separates ethically misaligned or misleading answers from calibrated, semantically valid responses. Given a query-answer embedding $h_{\text{anchor}}$, we identify a positive sample $h_{\text{pos}}$ (semantically similar and ethically aligned) and a negative sample $h_{\text{neg}}$ (incorrect, overconfident, or potentially misleading). The contrastive loss is defined as according to Equation (3).
$\mathcal{L}_{\text{contrast}}=$ $\max\Big{(}0,\ m-\text{sim}(h_{\text{anchor}},h_{\text{pos}})$
@@
where $\text{sim}(\cdot)$ denotes cosine similarity, and $m$ is a margin hyperparameter. We use a projection head atop the DocVQA encoder to map answer embeddings into a low-dimensional calibrated honesty space. Where, positive pairs are identified using a combination of Word Mover’s Distance (WMD) and agreement with ground truth as shown in Equations (4), and (5), where $\delta$ is a tunable similarity threshold.
@@
$\hat{y}\in\mathcal{A}_{\text{aligned}}\Longrightarrow$ semantically valid
@@
Whereas, negative samples are drawn from high-confidence as shown in Equation (6).
$\hat{y}_{\text{neg}}:$ $\mathbb{1}[\hat{y}_{\text{neg}}\neq y^{*}]$
@@ 3.4 Training Module
Here, $\lambda_{1}$ and $\lambda_{2}$ control the relative weight of each loss term. Training is conducted end-to-end using batches sampled from standard DocVQA datasets, where each sample includes contrastive triplets and confidence-aware labels. The projection head is fine-tuned during training, while the base DocVQA encoder remains frozen. At inference, $\mathcal{C}(D,Q)$ and $\mathcal{U}(D,Q)$ may be used to suppress or abstain from answering under high uncertainty. Note: In this work, we define ethical calibration as the act of reducing confidently incorrect answers, especially under ambiguity or lack of sufficient visual-textual grounding. HonestVQA is not a moral arbiter but a mechanism for promoting caution and transparency in DocVQA behavior.
@@ 4.2 Evaluation Metrics
To comprehensively assess the performance and ethical alignment of our framework, we employ standard accuracy metrics alongside two novel measures designed to evaluate calibration and honesty. First, we report conventional Accuracy and F1 scores (i.e. macro) to quantify answer correctness. To capture the alignment between model confidence and actual correctness, we introduce the Honesty Score (H-Score), which penalizes overconfident incorrect predictions while rewarding calibrated confidence on correct answers. Additionally, the Ethical Confidence Index (ECI) evaluates the model’s ability to appropriately express uncertainty, especially under ambiguous or insufficient information. Note: In the tables, color coding is used to indicate the performance of the models: Blue represents the best performance, Green indicates the second-best performance, and Yellow denotes the third-best performance. Whereas, ↑ indicates that a high value is preferable, while ↓ indicates that a low value is preferable.
@@ 4.2.1 Theoretical Guarantees for Evaluation Metrics
In this section, we provide formal lemmas and proofs to establish the theoretical soundness of the proposed Honesty Score (H-Score) and Ethical Confidence Index (ECI), which measure the alignment between model confidence, accuracy, and ethical transparency in DocVQA.
@@
Let $C(D,Q)$ be the confidence score output by a DocVQA model for a given document-question pair $(D,Q)$, and let $A(D,Q)\in\{0,1\}$ be the corresponding accuracy indicator, where $1$ denotes a correct answer and 00 an incorrect one. Assume that $C(D,Q)$ is bounded in $[0,1]$. Then, define the Honesty Score $\mathrm{H}$ as according to Equation (8).
@@
By definition, calibration error measures the absolute difference between predicted confidence and true correctness. Since $A(D,Q)$ is binary, the expectation of $|C-A|$ captures the average misalignment. Rearranging, $\mathrm{H}=1-\mathbb{E}[|C-A|]$. Because $|C-A|\in[0,1]$, $\mathrm{H}\in[0,1]$ and is maximized when confidence perfectly matches accuracy. Hence, $\mathrm{H}$ is a valid measure of calibration that upper-bounds expected miscalibration. ∎
@@
Table 1: Answer correctness comparison on SpDocVQA, InfographicsVQA, and SROIE datasets
@@
Table 2: Calibration metrics on SpDocVQA, InfographicsVQA, and SROIE datasets
@@ 4.3 Hyperparameters
The HonestVQA framework employs several key hyperparameters to balance confidence calibration and contrastive learning effectively. We set the confidence penalty weight $\alpha$ to 1.0 and the cross-entropy weight $\beta$ to 0.5 to emphasize penalizing overconfident incorrect predictions while maintaining prediction accuracy. The contrastive margin $m$ is fixed at 0.3 to enforce a moderate separation between positive and negative embeddings. The alignment and contrastive losses are weighted by $\lambda_{1}=1.0$ and $\lambda_{2}=0.7$, respectively, reflecting a slightly stronger emphasis on alignment. For sample selection in the contrastive module, the WMD threshold $\delta$ is set to 0.4, while the confidence and entropy thresholds, $\tau_{1}$ and $\tau_{2}$, are chosen as 0.8 and 0.5, respectively, to effectively identify semantically valid positive samples and high-confidence misleading negatives. Note: We will make the code publicly available post-acceptance.
5 Experimental Analysis
5.1 Comparison with Baselines
We evaluate the effectiveness of HonestVQA by measuring both standard answer correctness metrics and calibration-specific metrics to provide a comprehensive assessment of the model’s performance. Specifically, we compare the base DocVQA models—LayoutLMv3 Fujitake (2024), UDOP Wang et al. (2023a), and DONUT Li et al. (2024)—with their corresponding versions enhanced by the HonestVQA calibration framework. Table 1 presents these results on the three datasets. It is evident that HonestVQA consistently improves accuracy by approximately 3 to 4 percentage points and macro F1-score by nearly 4 points across all base models. For instance, LayoutLMv3 Fujitake (2024) improves from 72.3% to 75.9% in accuracy and from 68.5% to 72.8% in macro F1-score. Similar trends hold for UDOP Wang et al. (2023a) and DONUT Li et al. (2024) models, underscoring the robustness of our approach in enhancing answer correctness. Whereas, Table 2 displays these calibration-specific results for the same set of models and dataset. Notably, the base models exhibit relatively high H-Score and ECI values, indicating frequent instances of unjustified overconfidence. Incorporation of HonestVQA substantially lowers these values, with H-Score decreasing by over 35% and ECI by nearly 40% on average. For instance, LayoutLMv3’s Fujitake (2024) H-Score drops from 0.185 to 0.113, and ECI decreases from 0.210 to 0.132 after calibration. This demonstrates that HonestVQA effectively mitigates the risk of misleading the user by suppressing confident but incorrect answers.
@@ 5.1 Comparison with Baselines
Table 3: Cross-domain hallucination detection performance of HonestVQA
@@
Table 4: Ablation results across three datasets showing the impact of disabling individual components of HonestVQA

### v2

@@ Abstract
Document Visual Question Answering (DocVQA) models often produce overconfident or ethically misaligned responses, especially under uncertainty. Existing models like LayoutLMv3, UDOP, and DONUT focus on accuracy but lack ethical calibration. We propose HonestVQA, a model-agnostic, self-supervised framework that aligns model confidence with correctness using weighted loss and contrastive learning. We introduce two new metrics—Honesty [^1] In our work, honesty is operationalized as the alignment between model confidence and correctness, with the goal of reducing confidently wrong answers. Score (H-Score) and Ethical Confidence Index (ECI)—to evaluate ethical alignment. HonestVQA improves accuracy and F1 by up to 4.3% across SpDocVQA, InfographicsVQA, and SROIE datasets, while reducing overconfidence. It also generalizes well across domains, achieving 78.9% accuracy and 76.1% F1-score.
@@
Sahil Tripathi1, Md Tabrez Nafis [^1] footnotemark: 1, Imran Hussain1, Jiechao Gao2thanks: Corresponding Author: jiechao@stanford.edu, tabrez.nafis@gmail.com 1Jamia Hamdard, New Delhi, India 2Center for SDGC, Stanford University, California, USA
@@ 1 Introduction
Document Visual Question Answering (DocVQA) has emerged as a key challenge in multimodal AI Wang et al. (2025), enabling models to answer questions based on visual and textual content in documents such as invoices, forms, contracts, and academic papers. These models are widely deployed in enterprise automation Jiang et al. (2024), legal analysis Liu et al. (2024), and assistive technologies Zeng et al. (2025). However, despite their growing utility, DocVQA models often lack ethical transparency—frequently returning confidently incorrect answers to ambiguous, adversarial, or under-specified queries. For instance, a model may assert the total invoice amount with high confidence even when the relevant table is partially occluded, or confidently misinterpret a scanned signature line as a date. Such failures can propagate serious downstream consequences, including legal misinterpretation, misinformation, or financial misjudgment.
However, the crux of the problem lies in the inability of existing DocVQA models to communicate uncertainty in a calibrated, ethically responsible manner. While State-of-the-Art (SOTA) models such as LayoutLMv3222https://huggingface.co/microsoft/layoutlmv3-base Fujitake (2024), UDOP [^3] https://huggingface.co/microsoft/udop-large Wang et al. (2023a), and DONUT [^4] https://huggingface.co/naver-clova-ix/donut-base Li et al. (2024) focus on improving accuracy through sophisticated architecture and pretraining strategies, they fall short in aligning model confidence with actual knowledge. LayoutLMv3 Fujitake (2024) tends to prioritize exact answers over conveying doubt, UDOP Wang et al. (2023a) frequently errs on the side of over-caution without actionable explanations, and DONUT Li et al. (2024) offers no uncertainty estimation at all—leading to ethically untrustworthy behavior in ambiguous scenarios. Therefore, recent advances in AI alignment research have emphasized the importance of ethical calibration Rao et al. (2023), including honesty Yang et al. (2024), confidence-awareness Stangel et al. (2025), and transparent failure modes Stewart et al. (2023). However, these insights have yet to be meaningfully integrated into DocVQA models.
To address these critical gaps, we propose HonestVQA, a self-supervised framework that calibrates model confidence to reflect its underlying knowledge and ethical responsibility. Our approach is model-agnostic and integrates three key components: (1) uncertainty quantification to identify knowledge gaps, (2) confidence-accuracy alignment through weighted loss optimization, and (3) contrastive learning to enforce ethical response boundaries in ambiguous contexts. We also introduce two novel evaluation metrics: i) Honesty Score (H-Score), which captures the alignment between confidence and correctness, and ii) Ethical Confidence Index (ECI), which evaluates whether high-confidence answers are ethically warranted.
@@ 2 Related Work
Recent research has increasingly focused on improving the reliability and interpretability of AI models, especially in high-stakes domains. While core DocVQA models like have been discussed in Section 1, here we focus on complementary areas that our framework draws from—confidence calibration, ethical modeling, and contrastive learning. Confidence calibration models such as temperature scaling Xie et al. (2024) and label smoothing Müller et al. (2019) aim to align predicted probabilities with empirical accuracies. However, these models are typically post-hoc and task-agnostic, often failing to generalize in multimodal settings. Selective prediction frameworks such as Chen et al. (2023) allow models to abstain from uncertain answers, but they usually rely on fixed thresholds and lack principled mechanisms to model epistemic uncertainty in visually grounded tasks like DocVQA. However, in the area of ethical and honest AI, efforts such as instruction tuning for alignment Zhang et al. (2023) and calibrated language modeling Zhu et al. (2023) emphasize epistemic humility—training models to express uncertainty when appropriate. However, these researches are primarily developed for language-only models and remain underexplored in multimodal tasks involving structured visual data. Whereas, contrastive learning has shown strong performance in aligning multimodal representations, with frameworks like CLIP Gao et al. (2024) and ALIGN Wang et al. (2023b) leveraging contrastive objectives for image-text alignment. While effective at learning generalizable embeddings, such models are not designed to enforce ethical boundaries or distinguish between honest and overconfident outputs in ambiguous scenarios.
@@
[…]
@@
[…]
@@ 3 Methodology
As discussed earlier, HonestVQA is a model-agnostic calibration framework designed to enhance ethical transparency in DocVQA models. It operates as a wrapper around pretrained DocVQA models (in our work, we evaluate our framework on top of pretrained models such as LayoutLMv3 Fujitake (2024), UDOP Wang et al. (2023a), and DONUT Li et al. (2024) to demonstrate its generalizability), injecting uncertainty-aware alignment and contrastive reasoning to reduce overconfident yet incorrect outputs. The broader process of the HonestVQA is illustrated in Algorithm 1.
@@ 3.3 Contrastive Ethical Enforcement Module
To further refine the model’s response space under ambiguity, we introduce a contrastive loss that structurally separates ethically misaligned or misleading answers from calibrated, semantically valid responses. Given a query-answer embedding $h_{\text{anchor}}$, we identify a positive sample $h_{\text{pos}}$ (semantically similar and ethically aligned) and a negative sample $h_{\text{neg}}$ (incorrect, overconfident, or potentially misleading). The contrastive loss is defined as according to Equation (4), where $\text{sim}(\cdot)$ denotes cosine similarity, and $m$ is a margin hyperparameter.
$\mathcal{L}_{\text{contrast}}=\$$\max\Big(0,\ m-\text{sim}(h_{\text{anchor}},h_{\text{pos}})$
@@
We use a projection head atop the DocVQA encoder to map answer embeddings into a low-dimensional calibrated honesty space. Where, positive pairs are identified using a combination of Word Mover’s Distance (WMD) and agreement with ground truth as shown in Equations (5), and (6), where $\delta$ is a tunable similarity threshold.
@@
$\hat{y}\in\mathcal{A}_{\text{aligned}}\Longrightarrow\$semantically valid
@@
Whereas, negative samples are drawn from high-confidence as shown in Equation (7), where $\tau_{1}$ and $\tau_{2}$ are confidence and entropy thresholds, respectively.
$\hat{y}_{\text{neg}}:\$$\mathbb{1}[\hat{y}_{\text{neg}}\neq y^{*}]$
@@ 3.4 Training Module
Here, $\lambda_{1}$ and $\lambda_{2}$ control the relative weight of each loss term. Training is conducted end-to-end using batches sampled from standard DocVQA datasets, where each sample includes contrastive triplets and confidence-aware labels. The projection head is fine-tuned during training, while the base DocVQA encoder remains frozen. At inference, $\mathcal{C}(D,Q)$ and $\mathcal{U}(D,Q)$ may be used to suppress or abstain from answering under high uncertainty. Note: In this work, we define ethical calibration as the act of reducing confidently incorrect answers, especially under ambiguity or lack of sufficient visual-textual grounding. HonestVQA is not a moral arbiter but a mechanism for promoting caution and transparency in DocVQA behavior. Furthermore, our framework is fully multimodal. HonestVQA operates as a wrapper on LayoutLMv3, UDOP, and DONUT—all of which jointly model text + layout + vision. The uncertainty and contrastive modules explicitly leverage multimodal embeddings (e.g., attention-grounded text regions). This is confirmed by our IoU multimodal consistency evaluation (see Table 6), where HonestVQA achieves significantly higher alignment between visual attention and textual grounding compared to baselines.
@@ 4.2 Evaluation Metrics
To comprehensively assess the performance and ethical alignment of our framework, we employ standard accuracy metrics alongside two novel measures designed to evaluate calibration and honesty. First, we report conventional Accuracy and F1 scores (i.e. macro) to quantify answer correctness. To capture the alignment between model confidence and actual correctness, we introduce the H-Score, which penalizes overconfident incorrect predictions while rewarding calibrated confidence on correct answers. Additionally, the ECI evaluates the model’s ability to appropriately express uncertainty, especially under ambiguous or insufficient information. Note: In the tables, Green indicate the best-performing scores. ↑ indicates that a high value is preferable, while ↓ indicates that a low value is preferable.
@@ 4.2.1 Theoretical Guarantees for Evaluation Metrics
In this section, we provide formal lemmas and proofs to establish the theoretical soundness of the proposed H-Score and ECI, which measure the alignment between model confidence, accuracy, and ethical transparency in DocVQA.
@@
Let $C(D,Q)$ be the confidence score output by a DocVQA model for a given document-question pair $(D,Q)$, and let $A(D,Q)\in\{0,1\}$ be the corresponding accuracy indicator, where $1$ denotes a correct answer and 0 an incorrect one. Assume that $C(D,Q)$ is bounded in $[0,1]$. Then, define the Honesty Score $\mathrm{H}$ as according to Equation (9).
@@
By definition, calibration error measures the absolute difference between predicted confidence and true correctness. Since $A(D,Q)$ is binary, the expectation of |CA||C-A| captures the average misalignment. Rearranging, $\mathrm{H}=1-\mathbb{E}[|C-A|]$. Because $|C-A|\in[0,1]$, $\mathrm{H}\in[0,1]$ and is maximized when confidence perfectly matches accuracy. Hence, $\mathrm{H}$ is a valid measure of calibration that upper-bounds expected miscalibration. ∎
@@
Table 1: Performance comparison of base DocVQA models and their HonestVQA-enhanced versions across SpDocVQA, InfographicsVQA, and SROIE datasets, showing improvements in both Accuracy and Macro F1 scores.
@@
Table 2: Calibration performance of base DocVQA models and their HonestVQA-enhanced versions on SpDocVQA, InfographicsVQA, and SROIE datasets, measured by H-Score and ECI. Lower values indicate better model calibration and reduced overconfidence.
@@ 4.3 Hyperparameters
The HonestVQA framework employs several key hyperparameters to balance confidence calibration and contrastive learning effectively. We set the confidence penalty weight $\alpha$ to 1.0 and the cross-entropy weight $\beta$ to 0.5 to emphasize penalizing overconfident incorrect predictions while maintaining prediction accuracy. The contrastive margin $m$ is fixed at 0.3 to enforce a moderate separation between positive and negative embeddings. The alignment and contrastive losses are weighted by $\lambda_{1}=1.0$ and $\lambda_{2}=0.7$, respectively, reflecting a slightly stronger emphasis on alignment. For sample selection in the contrastive module, the WMD threshold $\delta$ is set to 0.4, while the confidence and entropy thresholds, $\tau_{1}$ and $\tau_{2}$, are chosen as 0.8 and 0.5, respectively, to effectively identify semantically valid positive samples and high-confidence misleading negatives. Note: At inference, the model does not access ground-truth answers. Instead, it relies on the learned calibration from training–the confidence score $C(D,Q)$ and entropy $U(D,Q)$ are computed from the predicted distribution to identify uncertain or overconfident answers. The model can then suppress or abstain from outputs under high uncertainty, while the contrastive embeddings help separate ethically aligned responses from potentially misleading ones.
@@
Table 3: Cross-domain performance of HonestVQA, showing Accuracy and Macro F1 when trained on one dataset and evaluated on another. Results indicate robust generalization across SpDocVQA, InfographicsVQA, and SROIE datasets.
@@
Table 4: Ablation study of HonestVQA showing the impact of removing alignment or contrastive loss on Accuracy and H-Score across SpDocVQA, InfographicsVQA, and SROIE. The full model consistently achieves the best performance and calibration.
5 Experimental Analysis
5.1 Comparison with Baselines
We evaluate the effectiveness of HonestVQA by measuring both standard answer correctness metrics and calibration-specific metrics to provide a comprehensive assessment of the model’s performance. Specifically, we compare the base DocVQA models—LayoutLMv3 Fujitake (2024), UDOP Wang et al. (2023a), and DONUT Li et al. (2024)—with their corresponding versions enhanced by the HonestVQA calibration framework. Table 1 presents these results on the three datasets. It is evident that HonestVQA consistently improves accuracy by approximately 3% to 4% and macro F1-score by nearly 4% across all base models. For instance, LayoutLMv3 Fujitake (2024) improves from 72.3% to 75.9% in accuracy and from 68.5% to 72.8% in macro F1-score. Similar trends hold for UDOP Wang et al. (2023a) and DONUT Li et al. (2024) models, underscoring the robustness of our approach in enhancing answer correctness.
Whereas, Table 2 displays these calibration-specific results for the same set of models and dataset. Notably, the base models exhibit relatively high H-Score and ECI values, indicating frequent instances of unjustified overconfidence. Incorporation of HonestVQA substantially lowers these values, with H-Score decreasing by over 35% and ECI by nearly 40% on average. For instance, LayoutLMv3’s Fujitake (2024) H-Score drops from 0.185 to 0.113, and ECI decreases from 0.210 to 0.132 after calibration. This demonstrates that HonestVQA effectively mitigates the risk of misleading the user by suppressing confident but incorrect answers.